OpenStudy (anonymous):

If the rotation matrix in the rotation equation is matrix[x_y]=matrix[-√(2)/2,-√(2)/2,_√(2)/2,-√(2)/2,]*matrix[x'_y']. What angle of rotation is being used?

OpenStudy (anonymous):

I just wrote a huge piece and this site kiboshed what I wrote. Did you get anything for this question?

OpenStudy (anonymous):

Basically, identify the entries in your matrix with each of the trig. functions that describe a rotation matrix in two dimensions. You should have then,\[\sin \theta = -\frac{1}{\sqrt{2}}\]and \[\cos \theta =-\frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

In one rotation, \[\sin \theta = -\frac{1}{\sqrt{2}} \rightarrow \theta =225^o, 315^o\]

OpenStudy (anonymous):

\[\cos \theta = -\frac{1}{\sqrt{2}} \rightarrow \theta = 135^o, 225^o\]

OpenStudy (anonymous):

The only angle BOTH trig. functions have in common is 225 degrees. Unless I've misread your matrix entries, this is the angle of rotation.